I am trying to help my son with his prealgebra but it has been along time for me. Here is the inequality. -(x/5)<_0
It reads as x divided by 5, negative, is less than or equal to 0.

Found 2 solutions by Alan3354, rfer:
Answer by Alan3354(69443)

-(x/5)<=0
-x<=0
x>=0 (note change of polarity)
x is greater than or equal to zero

Answer by rfer(16322)

-x/5<=0
multiply both sides by 5
since 5*0=0
-x<=0
to clear the - from x
multiply both sides by -1
when multiplying in inequality by a - the <= gets reversed
x>=0